please help!! okay how do you solve this proportion?
(3x-2)/4=(x+4)/2

Answer by wuwei96815(245)

(3x-2)/4=(x+4)/2
Multiply both sides by 4.
(3x-2)=2(x+4)
Remvove parentheses.
3x-2=2x+8
Add -2x to both sides.
x-2=8
Add 2 to both sides.
x=10

Did you follow that? To check our answer go back and substitute 10 for x in the original equation.